Shift start with control access. RHRS magnets are ready; LHRS dipole is still down. Both VDC off. LD2 target.
16:13 RHRS Q3 is cycled. Didn't see Q2 history. So cycling it to 1400A and back.
16:31 LHRS Dipole recovered (miraclely)
16:41 Beam permit; Beam recover with Compton
17:13 Harp Scan:
17:15 Establishing New Beam Recovery Procedure (for BPM A = B = (0 , 0)mm) by Yi Qiang
18:00 Asked for 100uA beam for production
18:09 Experts finish the checking. Starting production run 13868/4917/25366
18:21 Beam off for some hall B tuning. Stop run.
20:47 Control Access for quick fixing on PVDIS DAQ narrow trigger
21:07 Control Access ended
21:33 Beam back. Starting tuning precedure @ new beam location = BPM A = B = (-0.7,0.0)mm.
22:16 Start Production Run : 11870/25373/4925
22:13 Start Production Run : 11871/25374/4926